### v1

@@ 7 Conclusions
To characterize the dense core population and star formation activities in the Central Molecular Zone (CMZ) of the Galaxy, we have carried out the Dual-band Unified Exploration of Three CMZ Clouds (DUET) survey of three representative clouds, the 20 km s-1 cloud, Sgr C, and Dust Ridge cloud e. DUET adopts cloud-wide mosaicked observations with ALMA, achieving a comparable resolution of 0.2–0.3″ (about 2,000 au at the distance of the CMZ) and a sky coverage of 8.3–10.4 arcmin2 at the 1.3 mm and 3 mm bands, respectively. In this first paper of the series, we provide an overview of the survey, release the continuum images, and report mono-band and dual-band source catalogs. Our results and discussions are summarized as follows.
Catalog description: using the getsf source extraction algorithm, we obtain mono-band catalogs (mbcat) of 563 sources at 1.3 mm, including 81 in cloud e, 224 in Sgr C, and 258 in the 20 km s-1 cloud, and 330 sources at 3 mm, including 39 in cloud e, 128 in Sgr C, and 163 in the 20 km s-1 cloud. Combining the two bands, we obtain dual-band catalogs (dbcat) of 450 sources, including 62 in cloud e, 182 in Sgr C, and 206 in the 20 km s-1 cloud. The two types of catalogs have their own advantages: mbcat is optimized for monochromatic completeness, while dbcat improves robustness and enables detections of faint and blended sources through dual-band cross-check.
@@
Low spectral indices and possible explanations: dbcat sources enable studies of the inter-band (3 mm–1.3 mm) spectral index (). To avoid the systematic bias from different ranges during the ALMA observations, we select the same ranges and re-clean the continuum images for both bands. With the same resolution and maximum recoverable scale, we find that % of the dbcat sources deviate from modified blackbody, being characterized either as the “low spectral indices” problem or as the “low brightness temperature” problem. To interpret the problems, we raise three hypotheses and discuss their possibilities:
@@
Class 0/I YSOs. Assuming optically thick disk sizes of 50, 100, and 200 au, the isothermal tracks are calculated based on the same dust model. The theoretical tracks can explain the observations. The hypothesis is further supported by recent long-baseline ALMA observations towards the CMZ clouds that have resolved candidates of envelopes or disks. The hypothesis is also reconciled with the fact that there are always a large number of beam-like protostellar objects regardless of the spatial resolutions of observations, because they have hardly been resolved.
@@
Existence of large (mm/cm-sized) grains. With a self-consistent dust radiation transfer model, we calculate isothermal tracks with various grain sizes from 10 m to 10 cm. We find that cm-sized grains can explain most of sources with low values, because non-negligible scattering flattens the frequency dependence of effective opacity and reduces at millimeter band. More importantly, with mm-sized grains, the “high- bump” where a set of detections with can also be explained.
@@
Contamination of free-free emission. Cross-matching with previous K-band JVLA observations, the low spectral indices of a small amount of sources can be corrected by subtracting free-free emission contamination. Our calculation shows that even if free-free emission is only 30 Jy at K-band, it can still reduce by a magnitude of 1 for a mJy source in our sample. However, the needed number of massive protostars (embedded UCHii regions) are infeasible for the IMF of star clusters (even considering a top-heavy IMF), whereas the free-free emission from low-mass YSOs are unconstrainable by current facilities.
@@ Appendix A ALMA 3 mm observation supplements
We present the log of the ALMA 3 mm observations in Table 3, and the spectral window in Fig. 11. The exemplar observed spectra is averaged from the densest part of the Sgr C cloud 666With 4″-radius aperture centered at 17h44m40.3s, 29d28m14.7s..
@@ Appendix B Source extraction procedure and catalogs
Two scale-related inputs of the getsf algorithm are 1) the minimum size and 2) the maximum size of sources. The minimum size of sources for the images of the two bands are defined by the full width at half maximum of the synthesized beams as listed in column (4) of Table 2. The maximum size is used to avoid false positive detections of extended structures from flat background emission. We visually estimate it to be 3–4 beam sizes, and then give a more conservative value of 5 beam sizes () when extracting sources. The corresponding physical scale, 0.05 pc or 10,000 au at the distance of CMZ, is close to the largest identified source size in previous surveys with similar resolution (e.g. Motte et al. 2018b; Pouteau et al. 2022; Xu et al. 2024b). As shown in Fig. 4, the maximum source size is 08, which is much smaller than the adopted maximum size. Therefore, the adopted maximum source size of will not lead to incompleteness of the source catalog.
Our getsf procedure is spitted into three rounds. In the first round, single-wavelength (monochromatic) separation of structural components of sources and filaments from their backgrounds is performed individually on the Band-3 and Band-6 continuum images without primary beam corrections. To avoid false detections at the image boundary, pixels with a primary beam response lower than 0.3 are masked and do not participate in source extraction. This run prepares flattened components of sources and filaments over selected wavebands.
The second round focuses on creating monochromatic core catalogs for both 1.3 mm and 3 mm. In this round, the detect and measure procedures are used to spatially decompose the flattened components, detect the positions of sources, and then measure the sources individually for the two bands. Consequently, two mono-band core catalogs (mbcat), which only contain detections at their individual bands, are obtained. The mbcats are shown in Table 4 and 5.
| ID | R.A. | Dec. | Sig | PA | Size | |||
| (J2000) | (J2000) | () | (deg) | (au) | (mJy beam-1) | (mJy) | ||
| 20 km s-1 | ||||||||
| Sgr C | ||||||||
| cloud e | ||||||||
| ID | R.A. | Dec. | Sig | PA | Size | |||
| (J2000) | (J2000) | () | (deg) | (au) | (mJy beam-1) | (mJy) | ||
| 20 km s-1 | ||||||||
| Sgr C | ||||||||
| cloud e | ||||||||
In the third round, the detect and measure procedures are restarted to detect cores and remeasure their properties in the two bands. Core identification considers both bands simultaneously, which means that a marginal detection in one band will be excluded if there is no significant detection in the other band. For cores that are detected in one band with high SNRs, they are included in the combined dual-band catalog and measured in the other band. The dual-band catalog (dbcat) is shown in Table 6.
@@ Appendix C Source deconvolution from beam
For interferometric data, the ellipticity of the beam cannot be ignored. The relative angular position between the observed source and the beam ellipses must be carefully considered, especially for unresolved and marginally resolved sources. As illustrated in Figure 5, there is a clear peak in the position angle distribution in the Sgr C 1.3 mm mbcat, which aligns with the beam position angle. This indicates that a significant number of sources are marginally resolved, with their measured shapes predominantly determined by the beam size. Proper deconvolution of these sources is crucial. To address this, we utilize covariance matrices for a more accurate solution.
@@
Solid deconvolution solution demands the matrix to be positive definite, i.e., the eigenvalues and should be positive. As such, the elliptical model of the source deconvolved from the beam is obtained as , , and . If the matrix is semi-definite (non-negative eigenvalues) or indefinite (negative eigenvalues), the real physical scenario is that a source is completely unresolved. In such a case, the source size is undetermined.
@@ Appendix D Spectral index approximation
At (sub)millimeter and centimeter wavelengths, is approximately expressed by (cf. Hildebrand 1983). To demonstrate it, we adopt the protostellar opacity model from (Ossenkopf & Henning 1994) and make an extrapolation for the 1.3 mm and 3 mm wavelengths. As shown in Figure 12, we consider the grain either with thin or thick ice mantle, which are respectively shown in orange and green data points. Since the structures of our interest are mostly dense cores ( cm-3), the density regime for the model is chosen to be or cm-3. These two regimes are marked with square and round line markers. In the zoom-in panel (0.1–1 THz), the four opacity models are linearly extrapolated by the five nearest data points (0.35–1.3 m). The extrapolated opacities are cm2 g-1 and cm2 g-1. The dust opacity spectral index , defined by,
@@ Appendix F Modified blackbody isothermal emission track
Taking advantage of the dual-band observations, one can break the degeneracy between temperature and surface density in the optically thick regime. For dual-band observations, modified blackbody (MBB) emission Eq. (2) writes as
@@ Appendix G Centimeter counterpart cross-match and free-free emission subtraction
In Sgr C, there are four radio compact sources H1–H4. H1 has a cm flux of 2.3 mJy, which can contribute 2 mJy at 3 mm. It is associated with a compact source SgrC-mb#55 that is only detected in 3 mm with a total flux of 3.6 mJy. We suggest that this source could be a UCHii region (with a size of 2,000 au) with more than 50% of its 3-mm flux from free-free emission, but without detectable dust emission down to 50 Jy at 1.3 mm. H2 is not associated with any sources. H3 is associated with two 3 mm bright sources (the red dots) SgrC-db#25 and SgrC-db#30. The millimeter spectral index of SgrC-db#30 is as flat as which is consistent with free-free emission spectrum, while SgrC-db#25 has of 0.07, suggesting that there is dust emission. The 3 mm integrated flux of the two sources are 3.1 mJy, while the expected integrated flux at 23 GHz would be 3.6 mJy if the 3 mm flux is from free-free emission, which is lower than the measured cm flux of H3, 5 mJy. However, H4 is more extended along the north-south direction, so we cannot determine the cm flux within the extent of SgrC-db#1. Therefore, we only subtract the peak flux of 1 mJy to make the correction. The spectral index changes from 3.3 to 3.5 for SgrC-db#1.

### v2

@@ 7 Conclusions
To characterize the dense core population and star formation activities in the CMZ of the Galaxy, we have carried out the Dual-band Unified Exploration of three CMZ Clouds (DUET) survey of three representative clouds, the 20 km s-1 cloud, Sgr C, and the dust ridge cloud e. DUET adopts cloud-wide mosaicked observations with ALMA, achieving a comparable resolution of 02–03 (about 2000 au at the distance of the CMZ) and a sky coverage of 8.3–10.4 arcmin2 at the 1.3 mm and 3 mm bands, respectively. In this first paper of the series, we provide an overview of the survey, release the continuum images, and report mono-band and dual-band source catalogs. Our results and discussions are summarized as follows.
Catalog description: Using the getsf source extraction algorithm, we obtain mono-band catalogs (mbcat) of 563 sources at 1.3 mm, including 81 in cloud e, 224 in Sgr C, and 258 in the 20 km s-1 cloud, and 330 sources at 3 mm, including 39 in cloud e, 128 in Sgr C, and 163 in the 20 km s-1 cloud. Combining the two bands, we obtain dual-band catalogs (dbcat) of 450 sources, including 62 in cloud e, 182 in Sgr C, and 206 in the 20 km s-1 cloud. The two types of catalogs have their own advantages: mbcat is optimized for monochromatic completeness, while dbcat improves robustness and enables detections of faint and blended sources through a dual-band cross-check.
@@
Low spectral indices and possible explanations: dbcat sources enable studies of the inter-band (3 mm–1.3 mm) spectral index (). To avoid the systematic bias from different ranges during the ALMA observations, we selected the same ranges and re-cleaned the continuum images for both bands. With the same resolution and maximum recoverable scale, we find that % of the dbcat sources deviate from MBB, being characterized either as the “low spectral indices” problem or as the “low brightness temperature” problem. To interpret the problems, we raise three hypotheses and discuss their possibilities.
@@
Class 0/I YSOs. Assuming optically thick disk sizes of 50, 100, and 200 au, the isothermal tracks were calculated based on the same dust model. The theoretical tracks can explain the observations. The hypothesis is further supported by recent long-baseline ALMA observations towards the CMZ clouds that have resolved candidates of envelopes or disks. The hypothesis is also supported by the fact that there are always a large number of beam-like protostellar objects regardless of the spatial resolutions of observations, because they have hardly been resolved.
@@
Existence of millimeter or centimeter-sized large grains. With a self-consistent dust radiation transfer model, we calculated isothermal tracks with various grain sizes from 10 m to 10 mm. We find that centimeter-sized grains can explain most of sources with low values, because non-negligible scattering flattens the frequency dependence of effective opacity and reduces at millimeter band. More importantly, there is a “high- bump” for which a set of detections with can also be explained.
@@
Contamination of free-free emission. Cross-matching with previous K-band JVLA observations, the low spectral indices of a small amount of sources can be corrected by subtracting free-free emission contamination. Our calculation shows that even if free-free emission is only 30 Jy in the K-band, it can still reduce by a magnitude of 1 for a milli-Jansky source in our sample. However, the number of massive protostars (embedded UCHii regions) needed is infeasible for the IMF of star clusters (even considering a top-heavy IMF), whereas the free-free emission from low-mass YSOs cannot be constrained by current facilities.
@@ Appendix A ALMA 3 mm observation supplements
We present the log of the ALMA 3 mm observations in Table 5, and the spectral window in Fig. 11. The exemplar observed spectra is averaged from the densest part of the Sgr C cloud 888With 4″-radius aperture centered at 17h44m40.3s, 29d28m14.7s..
@@ Appendix B Source extraction procedure and catalogs
Two scale-related inputs of the getsf algorithm are 1) the minimum size and 2) the maximum size of sources. The minimum size of sources for the images of the two bands are defined by the full width at half maximum of the synthesized beams as listed in column (4) of Table 2. The maximum size is used to avoid false positive detections of extended structures from flat background emission. We visually estimate it to be 3–4 beam sizes, and then give a more conservative value of 5 beam sizes () when extracting sources. The corresponding physical scale, 0.05 pc or 10,000 au at the distance of CMZ, is close to the largest identified source size in previous surveys with similar resolution (e.g., Motte et al. 2018b; Pouteau et al. 2022; Xu et al. 2024b). As is shown in Fig. 4, the maximum source size is 08, which is much smaller than the adopted maximum size. Therefore, the adopted maximum source size of will not lead to incompleteness of the source catalog.
Our getsf procedure is spitted into three rounds. In the first round, single-wavelength (monochromatic) separation of structural components of sources and filaments from their backgrounds is performed individually on the band 3 and band 6 continuum images without primary beam corrections. To avoid false detections at the image boundary, pixels with a primary beam response lower than 0.3 are masked and do not participate in source extraction. This run prepares flattened components of sources and filaments over selected wavebands.
The second round focuses on creating monochromatic core catalogs for both 1.3 mm and 3 mm. In this round, the detect and measure procedures are used to spatially decompose the flattened components, detect the positions of sources, and then measure the sources individually for the two bands. Consequently, two mono-band core catalogs (mbcat), which only contain detections at their individual bands, are obtained. The mbcats are shown in Table 3 and 4.
In the third round, the detect and measure procedures are restarted to detect cores and remeasure their properties in the two bands. Core identification considers both bands simultaneously, which means that a marginal detection in one band will be excluded if there is no significant detection in the other band. For cores that are detected in one band with high S/Ns, they are included in the combined dual-band catalog and measured in the other band. The dual-band catalog (dbcat) is shown in Table 6.
@@ Appendix C Source deconvolution from beam
For interferometric data, the ellipticity of the beam cannot be ignored. The relative angular position between the observed source and the beam ellipses must be carefully considered, especially for unresolved and marginally resolved sources. As is illustrated in Figure 5, there is a clear peak in the position angle distribution in the Sgr C 1.3 mm mbcat, which aligns with the beam position angle. This indicates that a significant number of sources are marginally resolved, with their measured shapes predominantly determined by the beam size. Proper deconvolution of these sources is crucial. To address this, we utilize covariance matrices for a more accurate solution.
@@
The solid deconvolution solution demands the matrix to be positive definite; that is, the eigenvalues, and , should be positive. As such, the elliptical model of the source deconvolved from the beam is obtained as , , and . If the matrix is semi-definite (non-negative eigenvalues) or indefinite (negative eigenvalues), the real physical scenario is that a source is completely unresolved. In such a case, the source size is undetermined.
@@ Appendix D Spectral index approximation
At (sub)millimeter and centimeter wavelengths, is approximately expressed by (cf. Hildebrand 1983). To demonstrate it, we adopt the protostellar opacity model from (Ossenkopf & Henning 1994) and make an extrapolation for the 1.3 mm and 3 mm wavelengths. As is shown in Figure 12, we consider the grain either with thin or thick ice mantle, which are respectively shown in orange and green data points. Since the structures of our interest are mostly dense cores ( cm-3), the density regime for the model is chosen to be or cm-3. These two regimes are marked with square and round line markers. In the zoom-in panel (0.1–1 THz), the four opacity models are linearly extrapolated by the five nearest data points (0.35–1.3 m). The extrapolated opacities are cm2 g-1 and cm2 g-1. The dust opacity spectral index , defined by,
@@ Appendix F Modified blackbody isothermal emission track
Taking advantage of the dual-band observations, one can break the degeneracy between temperature and surface density in the optically thick regime. For dual-band observations, MBB emission Eq. (2) reads as
@@ Appendix G Centimeter counterpart cross-match and free-free emission subtraction
In Sgr C, there are four radio compact sources H1–H4. H1 has a cm flux of 2.3 mJy, which can contribute 2 mJy at 3 mm. It is associated with a compact source SgrC-mb#55 that is only detected in 3 mm with a total flux of 3.6 mJy. We suggest that this source could be a UCHii region (with a size of 2000 au) with more than 50% of its 3-mm flux from free-free emission, but without detectable dust emission down to 50 Jy at 1.3 mm. H2 is not associated with any sources. H3 is associated with two 3 mm bright sources (the red dots) SgrC-db#25 and SgrC-db#30. The millimeter spectral index of SgrC-db#30 is as flat as which is consistent with free-free emission spectrum, while SgrC-db#25 has of 0.07, suggesting that there is dust emission. The 3 mm integrated flux of the two sources are 3.1 mJy, while the expected integrated flux at 23 GHz would be 3.6 mJy if the 3 mm flux is from free-free emission, which is lower than the measured cm flux of H3, 5 mJy. However, H4 is more extended along the north-south direction, so we cannot determine the cm flux within the extent of SgrC-db#1. Therefore, we only subtract the peak flux of 1 mJy to make the correction. The spectral index changes from 3.3 to 3.5 for SgrC-db#1.